# Adversarial Domain Adaptation for Cross-user Activity Recognition Using Diffusion-based Noise-centred Learning

### v1

@@ Abstract
Human Activity Recognition (HAR) plays a crucial role in various applications such as human-computer interaction and healthcare monitoring. However, challenges persist in HAR models due to the data distribution differences between training and real-world data distributions, particularly evident in cross-user scenarios. This paper introduces a novel framework, termed Diffusion-based Noise-centered Adversarial Learning Domain Adaptation (Diff-Noise-Adv-DA), designed to address these challenges by leveraging generative diffusion modeling and adversarial learning techniques. Traditional HAR models often struggle with the diversity of user behaviors and sensor data distributions. Diff-Noise-Adv-DA innovatively integrates the inherent noise within diffusion models, harnessing its latent information to enhance domain adaptation. Specifically, the framework transforms noise into a critical carrier of activity and domain class information, facilitating robust classification across different user domains. Experimental evaluations demonstrate the effectiveness of Diff-Noise-Adv-DA in improving HAR model performance across different users, surpassing traditional domain adaptation methods. The framework not only mitigates distribution mismatches but also enhances data quality through noise-based denoising techniques.
@@ I Introduction
Figure 1: The design of Diff-Noise-Adv-DA.
However, existing diffusion-based domain adaptation methods often neglect the crucial role of noise in improving domain adaptation, which can carry valuable information in both the forward diffusion process and the reverse diffusion denoising process. To address these overlooked aspects, we propose a novel approach that harnesses the potential of noise within diffusion models for domain adaptation, termed Diffusion-based Noise-centered Adversarial Learning Domain Adaptation (Diff-Noise-Adv-DA). Our method integrates noise-centered learning into the diffusion process, enabling the model to develop robust activity classification models that generalize effectively across different user domains by enhancing the utilization of noise. This method transforms the forward diffusion and reverse denoising processes into adversarial learning phases, aligning feature distributions between source and target domains and improving the model’s ability to adapt to individual behavioral differences.
@@
Thus, the forward diffusion and reverse denoising processes are designed as adversarial learning phases. In the forward phase, the objective is to effectively distinguish between user and activity classes, acting as a discriminator and ensuring the model can identify unique characteristics of different users and their activities. Conversely, the reverse phase aims to obscure the data distribution related to classification and users, functioning as a generator. This adversarial interaction helps the model robustly align feature distributions across domains, reducing discrepancies and enhancing generalization capabilities.
@@
    Novel Deep Domain Adaptation Framework: We propose a novel deep domain adaptation framework, Diff-Noise-Adv-DA, specifically designed for cross-user HAR. This framework integrates the concepts of adversarial learning and generative modeling in a unique manner to exploit the knowledge embedded in the diffusion model for better domain adaptation.
@@
    Noise-centered Adversarial Learning: We develop a unique noise-centered adversarial learning mechanism that transforms the noise in both forward and reverse diffusion processes into critical information carriers for classification tasks. By integrating activity and domain class information into the noise, we enable the model to extract and leverage embedded information, significantly enhancing the robustness and adaptability of HAR models across different user domains.
The paper is organized as follows: it starts with a review of related work in sensor-based cross-user HAR and state-of-the-art methods in transfer learning and domain adaptation. Next, it delves into the specifics of the Diff-Noise-Adv-DA method, detailing its design principles and emphasis on noise-centered adversarial domain adaptation. Following this, we describe our experimental setup and compare Diff-Noise-Adv-DA with existing methods, highlighting its effectiveness in leveraging noise information within diffusion models. The paper concludes with a discussion of our findings and suggestions for future research directions in this field. The Diff-Noise-Adv-DA implementation is available at https://github.com/helloworld1973/diffusion_models.
@@ II-A Sensor-based Cross-user Activity Recognition
Sensor-based Human Activity Recognition (HAR) has emerged as a key component in the field of ubiquitous computing, providing deep insights into human behaviors by analyzing sensor data and contextual clues. This technology employs a variety of sensor types, each suited to different contexts, and can be grouped into five primary categories: smartphones and wearables, ambient sensors, device-free sensing, vision-based systems, and miscellaneous sensor types, as indicated in several studies [2][14]. Our research specifically focuses on the use of wearable sensors for HAR.
In machine learning, the problem of HAR using sensors is viewed as a challenge in time series classification [15]. Several classification techniques have been suggested, including Bayesian networks, Support Vector Machines, and Hidden Markov Models (HMM). Additionally, deep learning models have shown impressive results across many tasks. In the context of HAR, these deep learning methods [16] can automatically learn to identify and extract relevant features from extensive datasets. However, these methods typically rely on the assumption that both training and testing data share the same distribution, conforming to the independent and identically distributed ($i.i.d.$) assumption [2]. Yet, this assumption is often violated in practical applications, where training and testing data sets usually differ significantly, falling into what is known as out-of-distribution ($o.o.d.$) scenarios [17]. Our study explores the challenges of sensor-based HAR in these $o.o.d.$ conditions.
@@ III-A Problem formulation
III-B An overview of the Diff-Noise-Adv-DA method
@@
Figure 2: The overview of the Diff-Noise-Adv-DA method.
@@
Moreover, both the forward diffusion and reverse denoising processes of Diff-Noise-Adv-DA method function as adversarial learning phases. In the forward diffusion phase, the goal is to differentiate between user and activity classes, acting as a discriminator to help the model recognize distinct characteristics of different users and their activities. In contrast, the reverse denoising phase aims to obscure the classification-related data distribution, acting as a generator. This adversarial setup helps the model better align feature distributions across different users, reducing discrepancies and improving generalization.
@@ III-C Forward Diffusion Process
In this section, we compare the origianl design of forward diffusion process of diffusion model to our novel design of Diff-Noise-Adv-DA.
@@ 1. Forward Diffusion Process Definition:
This equation defines the forward diffusion process with the inclusion of classification information. The distribution of $\mathbf{x}_{t}$ given $\mathbf{x}_{t-1}$, activity label $\mathbf{c}_{a}$, and user label $\mathbf{c}_{u}$ is a Gaussian distribution with a mean that combines the scaled previous data ${\alpha_{t}}\mathbf{x}_{t-1}$ and additional terms $\mathbf{c}_{a}\cdot\gamma_{a}+\mathbf{c}_{u}\cdot\gamma_{u}$ based on the classification labels, and variance $\beta_{t}\mathbf{I}$. $\mathbf{c}_{a}$ and $\mathbf{c}_{u}$ represent the activity and user labels, respectively. $\gamma_{a}$ and $\gamma_{u}$ are vectors that encode the influence of the activity and user labels on the data. The mean of the Gaussian distribution now incorporates these classification labels to guide the noising process. $\beta_{t}=1-\alpha_{t}$ determines the amount of the noise added.
@@ III-D Reverse Denoising Process
In this section, we compare the origianl design of reverse denoising process of diffusion model to our novel design of Diff-Noise-Adv-DA.
@@ 4. Reverse Noising Process with Adversarial Regularization:
In this way, the forward diffusion and reverse denoising processes cooperate together and function as adversarial learning. During the forward diffusion phase, the focus is on distinguishing between different user and activity classes, acting as a discriminator to help the model identify the unique features of various users and their activities. Conversely, the reverse denoising phase aims to obscure the data distribution related to classification, functioning as a generator. This adversarial framework helps the model better align feature distributions across different users, reducing discrepancies and enhancing its ability to generalize.
@@ IV Experiments
IV-A Datasets and data distribution analysis
We conducted extensive experiments on public HAR dataset PAMAP2, which offer a variety of activities and sensor configurations. For PAMAP2, we choos user 1, 2, 5and 8. because other users’ activities are not complete. there are 12 activities, i.e. lying, sitting, standing, walking, running, cycling, Nordic walking, ascending stairs, descending stairs, vacuum cleaning, ironing, rope jumping.
@@ IV-A Datasets and data distribution analysis
TABLE I: Comparison of the average feature distributions distance of Activities and the average Bootstrapped proportion of Activities between users on three datasets
@@
In Table I, we analyze the differences in data distribution from both a holistic user and dataset perspective. The ’average distance’ represents the average observed distances of all activities, while the ’average Bootstrapped proportion’ signifies the average Bootstrapped proportion of activities. Our findings indicate that the OPPT and PAMAP2 datasets exhibit more significant differences in feature distribution between activities compared to the DSADS dataset. This suggests a greater variation in activity features within OPPT and PAMAP2. Moreover, when examining sub-activity data distribution differences between users, we observe more pronounced disparities within the OPPT and DSADS datasets. This suggests that users in these datasets engage in sub-activities in more distinct ways compared to those in the PAMAP2 dataset. These distinct sub-activities reflect different data distributions, providing suitable cross-user datasets for our subsequent research endeavors.
IV-B Experimental Setup
Our experimental setup involves training the diffusion model with the proposed metrics on source user data and evaluating its performance on target user data. We measure the improvement in generalization using standard classification metrics such as accuracy, F1-score, and precision.
@@ IV-C Results and Discussion
The experimental analysis aims to evaluate the performance of various domain adaptation methods on the PAMAP2 dataset, focusing on user adaptation tasks. Specifically, we compare the methods CORAL, SOT, DANN, TrC, and Diff-Noise-Adv-DA across six tasks involving user-to-user adaptation.
The tasks involve domain adaptation from one user to another, denoted as $X\rightarrow Y$ where $X$ and $Y$ are user numbers. The performance metrics for each method are presented in terms of accuracy percentages for the following tasks:
  • $1\rightarrow 2$, $1\rightarrow 5$, $2\rightarrow 1$, $2\rightarrow 5$, $5\rightarrow 1$, $5\rightarrow 2$
  • lying, sitting, standing, walking, running, cycling, Nordic walking, ascending stairs, descending stairs, vacuum cleaning, ironing
  • 11 activities
The table below summarizes the performance of the five methods across the six tasks:
Method 1$\rightarrow$2 1$\rightarrow$5 2$\rightarrow$1 2$\rightarrow$5 5$\rightarrow$1 5$\rightarrow$2
CORAL 46.95 32.69 43.25 38.21 50.95 44.45
SOT 71.22 69.82 67.32 74.26 64.54 74.63
DANN 53.64 47.66 43.47 52.96 50.03 63.65
TrC 67.23 62.58 69.31 66.15 64.10 72.88
Diff-Noise-Adv-DA 88.01 81.95 83.86 85.83 83.24 91.05
TABLE II: Performance comparison of different methods on user adaptation tasks in the PAMAP2 dataset.
@@
From the experimental results, it is evident that Diff-Noise-Adv-DA significantly outperforms all other methods across all six tasks. Below is a detailed analysis of each method’s performance:
  • CORAL: This method consistently shows the lowest performance across all tasks, with accuracy ranging from 32.69% to 50.95%. This indicates that CORAL struggles to effectively adapt between users in the PAMAP2 dataset.
  • SOT: SOT demonstrates a marked improvement over CORAL, with accuracies ranging from 64.54% to 74.63%. While it performs well, it still falls short compared to the top-performing method.
  • DANN: DANN performs moderately, with accuracies between 43.47% and 63.65%. Although it shows better adaptation capabilities than CORAL, it does not reach the effectiveness of SOT or TrC.
  • TrC: TrC shows strong performance with accuracies ranging from 62.58% to 72.88%. It consistently outperforms CORAL and DANN but is still outpaced by Diff-Noise-Adv-DA.
  • Diff-Noise-Adv-DA: This method achieves exceptional results, with accuracies ranging from 81.95% to 91.05%. Diff-Noise-Adv-DA leads in every task, demonstrating its superior capability in domain adaptation for user-to-user scenarios. The substantial margin by which it surpasses other methods highlights its effectiveness and robustness.
Our results demonstrate that the proposed cross-user data distribution alignment techniques Diff-Noise-Adv-DA significantly improve the generalization performance of HAR models. The new design of diffusion models leads to notable enhancements in recognizing activities across different users.

### v2

@@ Abstract
Human Activity Recognition (HAR) plays a crucial role in various applications such as human-computer interaction and healthcare monitoring. However, challenges persist in HAR models due to the data distribution differences between training and real-world data distributions, particularly evident in cross-user scenarios. This paper introduces a novel framework, termed Diffusion-based Noise-centered Adversarial Learning Domain Adaptation (DNA-DA), designed to address these challenges by leveraging generative diffusion modeling and adversarial learning domain adaptation techniques. Traditional HAR models often struggle with the diversity of user behaviors and sensor data distributions. DNA-DA innovatively integrates the inherent noise within diffusion models, harnessing its latent information to enhance domain adaptation. Specifically, the framework transforms noise into a critical carrier of activity and domain class information, facilitating robust classification across different user domains. Experimental evaluations demonstrate the effectiveness of DNA-DA in improving HAR model performance across different users, surpassing traditional domain adaptation methods. The framework not only mitigates distribution mismatches but also enhances data quality through noise-based denoising techniques.
@@ I Introduction
Figure 1: The design of DNA-DA.
However, existing diffusion-based domain adaptation methods often neglect the crucial role of noise in improving domain adaptation, which can carry valuable information in both the forward diffusion process and the reverse diffusion denoising process. To address these overlooked aspects, we propose a novel approach that harnesses the potential of noise within diffusion models for domain adaptation, termed Diffusion-based Noise-centered Adversarial Learning Domain Adaptation (DNA-DA). Our method integrates noise-centered learning into the diffusion process, enabling the model to develop robust activity classification models that generalize effectively across different user domains by enhancing the utilization of noise. This method transforms the forward diffusion and reverse denoising processes into adversarial learning domain adaptation phases, aligning feature distributions between source and target domains and improving the model’s ability to adapt to individual behavioral differences.
@@
Thus, the forward diffusion and reverse denoising processes are designed as adversarial learning domain adaptation. In the forward phase, the objective is to effectively distinguish between user and activity classes, acting as a discriminator and ensuring the model can identify unique characteristics of different users and their activities. Conversely, the reverse phase aims to obscure the data distribution related to classification and users, functioning as a generator. This adversarial interaction helps the model robustly align feature distributions across domains, reducing discrepancies and enhancing generalization capabilities.
@@
    Novel Deep Domain Adaptation Framework: We propose a novel deep domain adaptation framework, DNA-DA, specifically designed for cross-user HAR. This framework integrates the concepts of adversarial learning domain adaptation and generative modeling in a unique manner to exploit the knowledge embedded in the diffusion model for better domain adaptation.
@@
    Noise-centered Adversarial Learning Domain Adaptation: We develop a unique noise-centered adversarial learning domain adaptation mechanism that transforms the noise in both forward and reverse diffusion processes into critical information carriers for classification tasks. By integrating activity and domain class information into the noise, we enable the model to extract and leverage embedded information, significantly enhancing the robustness and adaptability of HAR models across different user domains.
The paper is organized as follows: it starts with a review of related work in sensor-based cross-user HAR and state-of-the-art methods in transfer learning and domain adaptation. Next, it delves into the specifics of the DNA-DA method, detailing its design principles and emphasis on noise-centered adversarial domain adaptation. Following this, we describe our experimental setup and compare DNA-DA with existing methods, highlighting its effectiveness in leveraging noise information within diffusion models. The paper concludes with a discussion of our findings and suggestions for future research directions in this field. The DNA-DA implementation is available at https://github.com/helloworld1973/DNA-DA.
@@ II-A Sensor-based Cross-user Activity Recognition
Sensor-based HAR has emerged as a key component in the field of ubiquitous computing, providing deep insights into human behaviors by analyzing sensor data and contextual clues. This technology employs a variety of sensor types, each suited to different contexts, and can be grouped into five primary categories: smartphones and wearables, ambient sensors, device-free sensing, vision-based systems, and miscellaneous sensor types, as indicated in several studies [2][14]. Our research specifically focuses on the use of wearable sensors for HAR.
In machine learning, the problem of HAR using sensors is viewed as a challenge in time series classification [15]. Several classification techniques have been suggested, including Bayesian networks, Support Vector Machines, and Hidden Markov Models. Additionally, deep learning models have shown impressive results across many tasks. In the context of HAR, these deep learning methods [16] can automatically learn to identify and extract relevant features from extensive datasets. However, these methods typically rely on the assumption that both training and testing data share the same distribution, conforming to the independent and identically distributed ($i.i.d.$) assumption [2]. Yet, this assumption is often violated in practical applications, where training and testing data sets usually differ significantly, falling into what is known as out-of-distribution ($o.o.d.$) scenarios [17]. Our study explores the challenges of sensor-based HAR in these $o.o.d.$ conditions.
@@ III-A Problem formulation
III-B An overview of the DNA-DA method
@@
Figure 2: The overview of the DNA-DA method.
@@
Moreover, both the forward diffusion and reverse denoising processes of DNA-DA method function as adversarial learning phases. In the forward diffusion phase, the goal is to differentiate between user and activity classes, acting as a discriminator to help the model recognize distinct characteristics of different users and their activities. In contrast, the reverse denoising phase aims to obscure the classification-related data distribution, acting as a generator. This adversarial setup helps the model better align feature distributions across different users, reducing discrepancies and improving generalization.
@@ III-C Forward Diffusion Process
In this section, we compare the origianl design of forward diffusion process of diffusion model to our novel design of DNA-DA.
@@ 1. Forward Diffusion Process Definition:
This equation defines the forward diffusion process with the inclusion of classification information. The distribution of $\mathbf{x}_{t}$ given $\mathbf{x}_{t-1}$, activity label $\mathbf{c}_{a}$, and user label $\mathbf{c}_{u}$ is a Gaussian distribution with a mean that combines the scaled previous data ${\alpha_{t}}\mathbf{x}_{t-1}$ and additional terms $\mathbf{c}_{a}\cdot\gamma_{a}+\mathbf{c}_{u}\cdot\gamma_{u}$ based on the classification labels, and variance $\beta_{t}\mathbf{I}$. $\mathbf{c}_{a}$ and $\mathbf{c}_{u}$ represent the activity and user labels, respectively. $\gamma_{a}$ and $\gamma_{u}$ are hyperparameters that control the influence of the activity and user labels on the data. The mean of the Gaussian distribution now incorporates these classification labels to guide the noising process. $\beta_{t}=1-\alpha_{t}$ determines the amount of the noise added.
@@ III-D Reverse Denoising Process
In this section, we compare the origianl design of reverse denoising process of diffusion model to our novel design of DNA-DA.
@@ 4. Reverse Noising Process with Adversarial Regularization:
In this way, the forward diffusion and reverse denoising processes cooperate together and function as adversarial learning domain adaptation. During the forward diffusion phase, the focus is on distinguishing between different user and activity classes, acting as a discriminator to help the model identify the unique features of various users and their activities. Conversely, the reverse denoising phase aims to obscure the data distribution related to classification, functioning as a generator. This adversarial framework helps the model better align feature distributions across different users, reducing discrepancies and enhancing its ability to generalize.
@@ IV Experiments
In this section, we evaluate the performance of our DNA-DA method through a series of comprehensive experiments focused on cross-user HAR.
IV-A Datasets and Experimental Setup
For sensor-based HAR research, OPPT, PAMAP2, and DSADS are the most representative and commonly used. We selected these datasets to cover a broad range of daily and sports-related activities to validate our cross-user HAR approach. Details about the subjects and activities common across these datasets are shown in Table I. Our focus is on IMU sensor (i.e. accelerometer and gyroscope) data collected from the lower right arm, simulating a smartwatch-based scenario.
TABLE I: Information on the Sensor-based HAR Datasets.
Dataset Subjects Common Activities
OPPT S1, S2, S3 1 standing, 2 walking, 3 sitting, 4 lying
PAMAP2 1, 5, 6
1 lying, 2 sitting, 3 standing, 4 walking,
5 running, 6 cycling, 7 Nordic walking,
8 ascending stairs, 9 descending stairs,
10 vacuum cleaning, 11 ironing
DSADS 2, 4, 7
1 sitting, 2 standing, 3 lying on back,
4 lying on right, 5 ascending stairs,
6 descending stairs, 7 standing in elevator still,
8 moving around in elevator, 9 walking in parking lot,
10 walking on treadmill in flat, 11 walking on treadmill inclined positions,
12 running on treadmill in flat, 13 exercising on stepper,
14 exercising on cross trainer, 15 cycling on exercise bike in horizontal positions,
16 cycling on exercise bike in vertical positions, 17 rowing, 18 jumping,
19 playing basketball
The OPPORTUNITY (OPPT) dataset [32] records subjects performing daily activities with minimal restrictions, following only a general description of tasks. The PAMAP2 dataset [33] features a more structured protocol with over 10 hours of data per participant, detailing specific activities for each person. Lastly, the Daily and Sports Activities Data Set (DSADS) [34] involves participants performing activities in their own way, leading to high variability that reflects real-world conditions. Each activity is performed for 5 minutes per participant.
These datasets are selected based on increasing levels of complexity for domain adaptation: OPPT with 4 activities, PAMAP2 with 11 activities, and DSADS with 19 activities. Notably, DSADS includes closely related activities like different types of walking, making it more challenging to classify.
For our experiments, we use a sliding window approach to segment the data, a common technique in sensor-based HAR [35]. Each window spans 3 seconds with a 50% overlap, capturing temporal relationships. We have approximately 20,000 sensor samples per activity for OPPT and PAMAP2, and 7,500 samples per activity for DSADS.
In the training phase, we use labeled data from a source user and unlabeled data from a target user. The target user’s data is split into a validation set and a test set: the first half serves as the validation set for tuning hyperparameters, while the second half is used for testing [6]. This temporal partitioning simulates new data being introduced over time. We select hyperparameters based on the validation set to ensure the model generalizes well and does not overfit. Finally, we evaluate the model’s performance on the test set using classification accuracy, assessing how effectively the model can classify activities for the target user.
IV-B Data Distribution Analysis
@@
TABLE II: Comparison of the average feature distributions distance of Activities and the average Bootstrapped proportion of Activities between users on three datasets
@@
In Table II, we analyze the differences in data distribution from both a holistic user and dataset perspective. The ’average distance’ represents the average observed distances of all activities, while the ’average Bootstrapped proportion’ signifies the average Bootstrapped proportion of activities. Our findings indicate that the OPPT and PAMAP2 datasets exhibit more significant differences in feature distribution between activities compared to the DSADS dataset. This suggests a greater variation in activity features within OPPT and PAMAP2. Moreover, when examining sub-activity data distribution differences between users, we observe more pronounced disparities within the OPPT and DSADS datasets. This suggests that users in these datasets engage in sub-activities in more distinct ways compared to those in the PAMAP2 dataset. These distinct sub-activities reflect different data distributions, providing suitable cross-user datasets for our subsequent research endeavors.
@@ IV-C Results and Discussion
The experimental analysis aims to evaluate the performance of various domain adaptation methods on the three datasets, focusing on user adaptation tasks. Specifically, we compare six methods of CORAL, SOT, DANN, OTDA, Raincoat and SPA to our DNA-DA.
  • CORAL [37]: CORAL (Covariate Shift Adaptation) adjusts the covariance matrices of feature layers to improve feature alignment across different domains, helping features become more domain-agnostic.
  • SOT [38]: The SOT (Substructure Optimal Transport) method explores the internal structure of domains to enable precise mapping at a substructure level. It effectively balances general and specific mapping requirements.
  • DANN [28]: DANN (Domain-Adversarial Neural Network) uses adversarial training to make a neural network’s features indistinguishable between source and target domains. The approach includes a discriminator that learns to identify whether features come from the source or target domain, while the network learns to produce features that fool the discriminator.
  • OTDA [39]: OTDA (Optimal Transport Domain Adaptation) employs optimal transport theory to efficiently map one probability distribution onto another, helping align samples from different domains.
  • Raincoat [40]: Raincoat addresses domain adaptation in time series by managing both feature and label shifts. It aligns features and labels across domains by integrating time and frequency information, and corrects any discrepancies between domains.
  • SPA [41]: The SPA (Spectral Alignment) method focuses on both inter-domain transferability and intra-domain structure. SPA transforms the domain adaptation problem into graph-based primitives, using a coarse graph alignment approach combined with a spectral regularizer to align domain graphs in eigenspaces.
Parameter Value
Training Epochs 500
Learning Rate 0.001
$\lambda_{\text{act}}$,$\lambda_{\text{binary}}$,$\lambda_{\text{adv}}$,$\lambda_{\text{act-source}}$ 1.0
$\gamma_{a}$,$\gamma_{u}$ 1.0
TABLE III: Parameter settings of the DNA-DA Method.
The primary metric for assessment is the classification accuracy for the target user. For the proposed DNA-DA method, parameter settings are listed in Table III. The different coefficients of constraint loss can be set based on the importance of these constraints. In our experiments setting, we unify the coefficients to be one to simplified the hyperparameters tuning process. In a similar way, we also unify the coefficients that control the influence of the activity and user labels on the data during forward diffusion process to be one. Better experiment results can be achieved by more fine-grained tuning of these coefficients.
@@
From the experimental results, it is evident that DNA-DA significantly outperforms all other methods across all the tasks. Below is a detailed analysis of each method’s performance:
Figure 8: OPPT dataset classification results.
On the OPPT dataset (see Figure 8), DNA-DA stands out with exceptional performance, achieving the highest accuracy across all scenarios. DNA-DA’s accuracy peaks at 97.94% in the S2 → S1 scenario, showcasing its superior adaptability and robustness in cross-subject activity recognition. Raincoat and TrC also perform commendably, with Raincoat achieving its highest accuracy of 90.44% in the S2 → S1 scenario, and TrC showing a peak accuracy of 86.99% in the S1 → S2 scenario. In contrast, other methods like OTDA and SA exhibit more moderate performance, with accuracies ranging from mid-70% to mid-80%. CORAL and SOT generally show lower performance, with accuracies below 80%, indicating their limitations in effectively managing the variability and complexity of the OPPT dataset.
Figure 9: PAMAP2 dataset classification results.
For the PAMAP2 dataset (see Figure 9), DNA-DA again demonstrates its robustness, with accuracy consistently above 76% and peaking at 79.96% in the 5 → 1 scenario. This highlights DNA-DA’s effective adaptation to the diverse set of activities in the PAMAP2 dataset. Raincoat also shows strong performance, especially in the 5 → 1 scenario (78.26%) and the 1 → 6 scenario (75.11%). TrC achieves its highest accuracy of 79.96% in the 5 → 6 scenario, but its overall performance is less consistent compared to DNA-DA. Methods such as DANN and OTDA display respectable accuracy in the range of 60% to 70%, while CORAL and SOT perform less effectively, with accuracies below 70%. This indicates that CORAL and SOT struggle to adapt to the structured and varied conditions present in the PAMAP2 dataset.
Figure 10: DSADS dataset classification results.
In the DSADS dataset (see Figure 10), characterized by high variability in activities, DNA-DA emerges as the top performer with the highest accuracy across all scenarios, peaking at 56.88% in the 7 → 2 scenario. This performance underscores DNA-DA’s ability to handle the dataset’s diverse and real-world conditions effectively. Raincoat and TrC also show reasonable performance, with Raincoat achieving up to 56.82% and TrC reaching 56.21%. Conversely, methods like DANN, OTDA, and SA exhibit lower performance, with accuracy in the range of 40% to 50%, highlighting their challenges in adapting to the variability and complexity of the DSADS dataset.
Overall, DNA-DA consistently proves to be the most effective method across all datasets, demonstrating superior adaptability and robustness in handling cross-user HAR tasks. DNA-DA’s ability to maintain high performance across different datasets underscores its robustness and generalizability. In comparison, Raincoat and TrC also show strong performance but are less consistent, with Raincoat performing particularly well in certain scenarios and TrC exhibiting variable results. Methods such as CORAL, SOT, and DANN display mixed results and generally perform less effectively, suggesting the need for further refinement to enhance their performance in diverse HAR contexts. The trends observed across these datasets highlight the reliability of DNA-DA and indicate potential areas for improvement in other methods. The experiments results demonstrate that the proposed cross-user data distribution alignment techniques DNA-DA significantly improve the generalization performance of HAR models.